# SciRIFF: A Resource to Enhance Language Model Instruction-Following over Scientific Literature

### v2

@@ 3.1 Evaluation
Table 1: Evaluation tasks included in SciRIFF-Eval. “/” separators indicate two separate subtasks. We use GPT-3.5 Turbo as our LLM judge and evaluate similarity on a 1-5 scale; see Appendix D.
@@
BioASQ List QA Question over a set of paper excerpts List of answer entities Exact match F1
BioRED Biomedical research abstract Mentions of 6 biomedical entity types Exact match F1
DiSCoMaT LaTex excerpt containing a table Values of table entries BLEU score vs. reference values
Evidence Inference (EI) Clinical trial report abstract (intervention, comparator, outcome, effect, evidence) tuples String overlap approximate match F1
MultiCite (MC) Citation context List of expressed citation intents (7 categories) Exact match F1
MuP Machine learning paper full text Short summary useful for a peer review Similarity to reference summary, evaluated by LLM judge
Qasper Question over an NLP paper full text Answer to question / Paragraphs providing attribution Similarity to reference answer, evaluated by LLM judge / Token F1 vs. reference paragraphs
SciERC Computer science abstract Mentions of 6 computer science entity types Exact match F1
SciFact Claim to verify against a biomedical research abstract Fact-checking verdict / Evidentiary sentences Label F1 / Token F1 vs. reference evidence
@@ 4.1 Main Results
SciTülu 7B outperforms Tülu V2 most substantially on information extraction tasks (BioRED, DiSCoMaT, and SciERC) [^10] SciTülu does not improve on Evidence Inference, which is an IE task. This likely occurs because the task is very difficult (Appendix D) and evaluation is challenging (§6), making current evaluations unreliable. Improving modeling and evaluation for complex IE tasks represents important future work. and tasks requiring attribution (Qasper, SciFact), achieving a 51% average improvement. On the other hand, it performs worse on summarization (MuP) and question answering (BioASQ). This may occur because summarization and QA are well-covered by existing instruction-following resources, while tasks like information extraction and attribution have received less attention.
@@ Appendix D Evaluation details
The following pages show full input / output examples for all SciRIFF-Eval tasks, along with details on metric calculations. This information is also available on the project GitHub page.

### v3

@@ 3.1 Evaluation
Table 1: Evaluation tasks included in SciRIFF-Eval. “/” separators indicate two separate subtasks. We use GPT-3.5 Turbo as our LLM judge and evaluate similarity on a 1-5 scale; see Appendix D. “Rel. extraction” stands for “Relation extraction”.
@@
BioASQ List QA QA Question over a set of paper excerpts List of answer entities Exact match F1
BioRED IE (NER) Biomedical research abstract Mentions of 6 biomedical entity types Exact match F1
DiSCoMaT IE (Table extraction) LaTex excerpt containing a table Values of table entries BLEU score vs. reference values
Evidence Inference (EI) IE (Rel. extraction) Clinical trial report abstract (intervention, comparator, outcome, effect, evidence) tuples String overlap approximate match F1
MultiCite (MC) Classification Citation context List of expressed citation intents (7 categories) Exact match F1
MuP Summarization Machine learning paper full text Short summary useful for a peer review Similarity to reference summary, evaluated by LLM judge
Qasper QA Question over an NLP paper full text Answer to question / Paragraphs providing attribution Similarity to reference answer, evaluated by LLM judge / Token F1 vs. reference paragraphs
SciERC IE (Rel. extraction) Computer science abstract Mentions of 6 computer science entity types Exact match F1
SciFact Entailment Claim to verify against a biomedical research abstract Fact-checking verdict / Evidentiary sentences Label F1 / Token F1 vs. reference evidence
@@ 4.1 Main Results
SciTülu 7B outperforms Tülu V2 most substantially on information extraction tasks (BioRED, DiSCoMaT, and SciERC) [^10] SciTülu does not improve on Evidence Inference, which is an IE task. This likely occurs because the task is very difficult (Appendix D) and evaluation is challenging (§6), making current evaluations unreliable. Improving modeling and evaluation for complex IE tasks represents important future work. and tasks requiring attribution (Qasper, SciFact), achieving a 51% average improvement. On the other hand, it performs worse on summarization (MuP) and question answering (BioASQ). This may occur because summarization and QA are well-covered by existing instruction-following resources, while tasks like information extraction and attribution have received less attention. It may also be an instance of negative task transfer (Jang et al., 2023; Asai et al., 2022), where training on many tasks at once can damage performance on a subset of target tasks. Based on these observations, we encourage practitioners using SciRIFF to conduct ablations to determine which training tasks are most beneficial for their particular use case.
@@ 6 Conclusion and Future Work
As presented in §2.2 the distribution of domains in SciRIFF concentrates heavily on AI and Bio / Clinical medicine. This reflects the availability of existing resources for scientific NLP. Looking forward, we hope that researchers will release new instruction-following resources covering a wider range of scientific domains, and we encourage the community to contribute novel tasks and datasets to SciRIFF as they become available.
@@ Appendix D Evaluation details
The following pages show full input / output examples for all SciRIFF-Eval tasks, along with details on metric calculations. This information is also available on the project GitHub page. For tasks using an LLM judge, we found in preliminary experiments that the results of GPT-3.5 were similar to other proprietary LLMs like GPT-4 and Claude-2; we used GPT-3.5 in the interest of cost and efficiency.